OpenStudy (anonymous):

Adding/Subtracting Fractions Can someone walk me through how to solve: 2x+5/x^2-3x-10 + x+1/x+2 Factors of 10 that equal negative 3 = -5 & 2 soo 2x+5/(x-5)(x+2) + (x+1)/(x+2)

OpenStudy (skullpatrol):

Rewrite the second expression (x+1)/(x+2) with the common denominator of (x-5)(x+2)

OpenStudy (anonymous):

So multiply the right side with (x-5)?

OpenStudy (skullpatrol):

Yes. Both the numerator and the denominator

OpenStudy (anonymous):

What about the left side? Don't I also have to multiply 2x+5/(x-5)(x+2) by something?

OpenStudy (skullpatrol):

Nope.

OpenStudy (anonymous):

Alright so 2x+5/(x-5)(x+2) + (+1)(x-5)/(x+2)(x-5) The (x-5)s cancel out and leave me with 2x+5/(x+2) + (x+1)/(x+2)(x-5)

OpenStudy (skullpatrol):

2x+5/(x-5)(x+2) + (x+1)(x-5)/(x+2)(x-5) Now you have common denominator, you can write the it all together

OpenStudy (anonymous):

I seem to have an (x-5) on the lower right hand side that hasn't cancelled out.

OpenStudy (anonymous):

Ohhh my goof

OpenStudy (skullpatrol):

[2x+5+ (x+1)(x-5)]/(x+2)(x-5)

OpenStudy (anonymous):

Yea I cancelled out instead of combining

OpenStudy (skullpatrol):

Now multiply out the numerator :-)

OpenStudy (anonymous):

So if I foil out (x+1)(x-5) I get 2x^2-4x-5 Then when I combine that with my original I get 2x^2-2x/(x-5)(x+2)

OpenStudy (skullpatrol):

2x + 5 + x^2 -4x - 5 = ?

OpenStudy (skullpatrol):

There is no 2x^2.

OpenStudy (anonymous):

Yea again my miscalculation

OpenStudy (anonymous):

2x + 5 + x^2 -4x - 5 = -4x+2x=-2x -5+5=0 x^2 cant combine into anything soo 2x + 5 + x^2 -4x - 5 = x^2-2x

OpenStudy (anonymous):

So x^2-2x/(x+2)(x-5)

OpenStudy (skullpatrol):

Yes. Check that you copied the original question correctly?

OpenStudy (anonymous):

Yep everything is in order, Thanks Skull!

OpenStudy (skullpatrol):

Good work :-)

OpenStudy (anonymous):

Thank you for your patience :P

OpenStudy (skullpatrol):

Thanks for trying to learn.